# Constructions of Efficiently Implementable Boolean Functions with Provable Nonlinearity/Resiliency/Algebraic Immunity Trade-Offs

### v2

@@ Abstract
We describe several families of efficiently implementable Boolean functions achieving provable trade-offs between resiliency, nonlinearity, and algebraic immunity. In particular, the following statement holds for each of the function families that we propose. Given integers $m_{0}\geq 0$, $x_{0}\geq 1$, and $a_{0}\geq 1$, it is possible to construct an $n$-variable function which has resiliency at least $m_{0}$, linear bias (which is an equivalent method of expressing nonlinearity) at most $2^{-x_{0}}$ and algebraic immunity at least $a_{0}$; further, $n$ is linear in $m_{0}$, $x_{0}$ and $a_{0}$, and the function can be implemented using $O(n)$ 2-input gates, which is essentially optimal.
@@ 1 Introduction
Theorem (informal): Given integers $m_{0}\geq 0$, $x_{0}\geq 1$ and $a_{0}\geq 1$, it is possible to construct an $n$-variable function which has resiliency at least $m_{0}$, linear bias at most $2^{-x_{0}}$ and algebraic immunity at least $a_{0}$, where $n$ is linear in $m_{0},x_{0}$ and $a_{0}$, and the function can be implemented using $O(n)$ 2-input gates, which is asymptotically optimal in the sense that any function which is at least $m_{0}$-resilient, has linear bias at most $2^{-x_{0}}$, and algebraic immunity at least $a_{0}$ must require $\Omega(\max(m_{0},x_{0},a_{0}))$ 2-input gates.
@@ 2 Background and Preliminary Results
For a positive integer $n$, the covering radius bound on the nonlinearity of an $n$-variable function $f$ is the following: ${\sf nl}(f)\leq 2^{n-1}-\lfloor 2^{n/2-1}\rfloor$, and equivalently, ${\sf LB}(f)\geq\lfloor 2^{n/2-1}\rfloor/2^{n-1}$. The bound is achieved if and only if $f$ is bent. Let $\chi(n)=\lfloor 2^{n/2-1}\rfloor/2^{n-1}$. We say that $f$ has almost optimal linear bias if $\chi(n)\leq{\sf LB}(f)\leq 2\chi(n)$, i.e. if the linear bias of $f$ is at most two times the lower bound arising from the covering radius bound. If $n$ is even, and ${\sf nl}(f)=2^{n-1}-2^{n/2}$, then ${\sf LB}(f)=2^{-(n-2)/2}=2\chi(n)$, while if $n$ is odd, and ${\sf nl}(f)=2^{n-1}-2^{(n-1)/2}$, then ${\sf LB}(f)=2^{-(n-1)/2}<2\chi(n)$; in both cases the linear bias is almost optimal.
@@
Theorem 2 holds for any affine map $\psi$. From efficiency considerations, we will consider $\psi$ to be a bit permutation, i.e. there is a permutation $\rho$ of $\{1,\ldots,k\}$ such that for any $(x_{1},\ldots,x_{k})\in\mathbb{F}_{2}^{n}$, $\psi(x_{1},\ldots,x_{k})=(x_{\rho(1)},\ldots,x_{\rho(k)})$. Implementation of a bit permutation requires only the proper connection pattern, and does not require any gates.
@@ 3 Basic Constructions of $m$-Resilient Functions
The main advantage of the functions constructed using Corollary 1 is that these functions can be constructed using $O(n)$ gates, while all previous constructions are essentially based on discrete logarithm computation and require super-polynomial size circuits. To see this advantage in concrete terms, we consider the case of the 28-variable function (say $f_{28}$) reported in Table 2 of [13] which has algebraic immunity 14 and nonlinearity 134201460 (equivalently linear bias equal to about $2^{-13.01}$). According to the description of hardware implementation in Section 4.2 of [13], implementation of $f_{28}$ will require a look-up table of size $2^{28}$ along with other gates. Taking $n=57$ in Corollary 1, we obtain a function (say $f_{57}$) with algebraic immunity 14 and linear bias equal to $2^{-28}$ which can be implemented using 217 NAND, 28 XOR, 30 AND, and 1 OR gates (see [7] for the method of obtaining the gate count). Comparing $f_{28}$ with $f_{57}$, we see that both are balanced and have the same algebraic immunity, $f_{57}$ has a much lower linear bias, and can be implemented much more efficiently than $f_{28}$. From both security and efficiency points of view, $f_{57}$ will be much more preferable than $f_{28}$ to a designer.
@@
For even $n$, several papers [25, 28, 22, 29, 23] have proposed constructions of 1-resilient functions with optimal algebraic immunity $n/2$ and provable upper bounds on linear bias. To the best of our knowledge the last of such results is [23] which improves upon works prior to it by providing lower linear bias, and also guarantees the fast algebraic immunity to be at least n6n-6. The upper bound on the linear bias of $n$-variable functions constructed in [23] is $((1+(n/2)\ln 2)/\pi+(\pi+16)/32)2^{-n/2+1}+2^{-(n-1)}$, whereas the linear bias of the functions constructed in Corollary 2 is $2^{-(n-2)/2}$ which is almost optimal, the algebraic immunity is at least $\lceil(n-2)/4\rceil$, and hence the fast algebraic immunity is at least $1+\lceil(n-2)/4\rceil$. So for 1-resilient functions, the functions constructed in [23] and those in Corollary 2 represent two distinct trade-off points with respect to algebraic immunity, fast algebraic immunity and linear bias. From an implementation point of view, however, there is a major difference between the constructions in Corollary 2 and that in [23]. The functions constructed using Corollary 2 require a circuit size of $O(n)$ gates, while the construction in [23] is based on defining the support of the desired Boolean function using powers of a primitive element over the field $\mathbb{F}_{2^{n/2}}$; in particular, from the description of Construction 2 in [23] it appears that the only reasonable method of implementing the function is to use a truth table requiring $\Omega(2^{n})$ gates. We provide a concrete example to highlight the difference between the functions constructed using Corollary 2 and Theorem 6 of [23] (which is based on Construction 2 of [23]). Suppose $n=32$. Theorem 6 of [23] provides a 32-variable, 1-resilient function (say $g_{32}$) with algebraic immunity 16, fast algebraic immunity at least 26 and nonlinearity at least 2147192232 (equivalently, linear bias at most about $2^{-12.85}$) which can be implemented using around $2^{32}$ gates. In Corollary 2 if we take $n=102$, then we obtain a 102-variable, 1-resilient function (say $g_{102}$) with algebraic immunity at least 25, and hence fast algebraic immunity at least 26, and almost optimal linear bias equal to $2^{-50}$ which can be implemented using 406 NAND, 52 XOR, 52 AND, and 3 OR gates (see [7] for the method of obtaining the gate count). So $g_{102}$ and $g_{32}$ both are 1-resilient and have the same fast algebraic immunity, $g_{102}$ has much higher algebraic immunity and much lower linear bias than $g_{32}$; and $g_{102}$ can be implemented much more efficiently than $g_{32}$. From both security and efficiency points of view, $g_{102}$ will be much more preferable than $g_{28}$ to a designer.
@@
The function $f_{5}$ is a 5-variable, 1-resilient function having degree 3, algebraic immunity 2, nonlinearity 12 (and hence almost optimal linear bias equal to $2^{-2}$), and can be implemented using 7 XOR gates and 4 AND gates.
@@
This shows that $f_{5}$ is the concatenation of 4 linear functions each of which is non-degenerate on at least 2 variables. It follows that $f_{5}$ is 1-resilient. It is easy to see that the degree of $f_{5}$ is 3. Further, it is not difficult to verify that the distance of $f_{5}$ to any affine is one of the values 12, 16 or 20, and so the nonlinearity of $f_{5}$ is 12. Hence, $f_{5}$ has almost optimal linear bias equal to $2^{-2}$. Since the degree of $f_{5}$ is 3, its algebraic immunity is at most 3. It is easy to see that neither $f_{5}$ nor $1\oplus f_{5}$ has any annihilator of degree 1. The following function is an annihilator of $f_{5}$: $Z_{1}X_{1}\oplus Z_{1}+Z_{2}\oplus X_{2}\oplus Z_{2}\oplus Z_{3}\oplus X_{1}\oplus Z_{3}X_{2}\oplus X_{1}X_{2}\oplus 1$. So the algebraic immunity of $f_{5}$ is 2. The expression for $f_{5}$ given by (6) can be implemented using 7 XOR gates and 4 AND gates.
@@ 5 Achieving Provable Resiliency/Nonlinearity/Algebraic Immunity Trade-Offs
Any $m_{0}$-resilient function is non-degenerate on at least $m_{0}+1$ variables, any function with linear bias at most $2^{-x_{0}}$ is non-degenerate on at least $2x_{0}$ variables, and any function with algebraic immunity at least $a_{0}$ is non-degenerate on at least $2a_{0}-1$ variables. The first statement now follows from Proposition 1. To see the second statement, note that in all the four cases of Theorem 9, the lower bound on $n$ is linear in $m_{0},x_{0}$ and $a_{0}$.

### v3

@@ Abstract
We describe several families of efficiently implementable Boolean functions achieving provable trade-offs between resiliency, nonlinearity, and algebraic immunity. In particular, the following statement holds for each of the function families that we propose. Given integers $m_{0}\geq 0$, $x_{0}\geq 1$, and $a_{0}\geq 1$, it is possible to construct an $n$-variable function which has resiliency at least $m_{0}$, linear bias (which is an equivalent method of expressing nonlinearity) at most $2^{-x_{0}}$ and algebraic immunity at least $a_{0}$; further, $n$ is linear in $\max(m_{0},x_{0},a_{0})$, and the function can be implemented using $O(n)$ 2-input gates, which is essentially optimal.
@@ 1 Introduction
Theorem (informal): Given integers $m_{0}\geq 0$, $x_{0}\geq 1$ and $a_{0}\geq 1$, it is possible to construct an $n$-variable function which has resiliency at least $m_{0}$, linear bias at most $2^{-x_{0}}$ and algebraic immunity at least $a_{0}$, where $n$ is linear in $\max(m_{0},x_{0},a_{0})$, and the function can be implemented using $O(n)$ 2-input gates, which is asymptotically optimal in the sense that any function which is at least $m_{0}$-resilient, has linear bias at most $2^{-x_{0}}$, and algebraic immunity at least $a_{0}$ must require $\Omega(\max(m_{0},x_{0},a_{0}))$ 2-input gates.
@@ 2 Background and Preliminary Results
For a positive integer $n$, the covering radius bound (see Theorem 3 in Chapter 3 of [4]) on the nonlinearity of an $n$-variable function $f$ is the following: ${\sf nl}(f)\leq 2^{n-1}-\lfloor 2^{n/2-1}\rfloor$, and equivalently, ${\sf LB}(f)\geq\lfloor 2^{n/2-1}\rfloor/2^{n-1}$. The bound is achieved if and only if $f$ is bent. Let $\chi(n)=\lfloor 2^{n/2-1}\rfloor/2^{n-1}$. We say that $f$ has almost optimal linear bias if $\chi(n)\leq{\sf LB}(f)\leq 2\chi(n)$, i.e. if the linear bias of $f$ is at most two times the lower bound arising from the covering radius bound. If $n$ is even, and ${\sf nl}(f)=2^{n-1}-2^{n/2}$, then ${\sf LB}(f)=2^{-(n-2)/2}=2\chi(n)$, while if $n$ is odd, and ${\sf nl}(f)=2^{n-1}-2^{(n-1)/2}$, then ${\sf LB}(f)=2^{-(n-1)/2}<2\chi(n)$; in both cases the linear bias is almost optimal.
@@
Theorem 2 holds for any affine map $\psi$. From efficiency considerations, we will consider $\psi$ to be a bit permutation, i.e. there is a permutation $\rho$ of $\{1,\ldots,k\}$ such that for any $(x_{1},\ldots,x_{k})\in\mathbb{F}_{2}^{n}$, $\psi(x_{1},\ldots,x_{k})=(x_{\rho(1)},\ldots,x_{\rho(k)})$. Implementation of a bit permutation requires only the proper connection pattern, and does not require any gates. For the case where $\psi$ is a bit permutation, it was shown in [14] that the algebraic immunity of $(\psi,{\sf Maj})\mbox{-}{\sf MM}_{2k}$ is at most $1+2^{\lceil\log(n/4)\rceil}$.
@@ 3 Basic Constructions of $m$-Resilient Functions
The main advantage of the functions constructed using Corollary 1 is that these functions can be constructed using $O(n)$ gates, while all previous constructions are essentially based on discrete logarithm computation and require super-polynomial size circuits. To see this advantage in concrete terms, we consider the case of the 28-variable function (say $f_{28}$) reported in Table 2 of [13] which has algebraic immunity 14 and nonlinearity 134201460 (equivalently linear bias equal to about $2^{-13.01}$). According to the description of hardware implementation in Section 4.2 of [13], implementation of $f_{28}$ will require a look-up table of size $2^{28}$ along with other gates. Taking $n=57$ in Corollary 1, we obtain a function (say $f_{57}$) with algebraic immunity 14 and linear bias equal to $2^{-28}$ which can be implemented using 217 NAND, 28 XOR, 30 AND, and 1 OR gates (see [7] for the method of obtaining the gate count). Comparing $f_{28}$ with $f_{57}$, we see that both are balanced and have the same algebraic immunity, $f_{57}$ has a much lower linear bias, and can be implemented much more efficiently than $f_{28}$. From both security and efficiency points of view, to a designer of cryptosystems $f_{57}$ will be much more preferable than $f_{28}$.
@@
For even $n$, several papers [26, 29, 23, 30, 24] have proposed constructions of 1-resilient functions with optimal algebraic immunity $n/2$ and provable upper bounds on linear bias. To the best of our knowledge the last of such results is [24] which improves upon works prior to it by providing lower linear bias, and also guarantees the fast algebraic immunity to be at least n6n-6. The upper bound on the linear bias of $n$-variable functions constructed in [24] is $((1+(n/2)\ln 2)/\pi+(\pi+16)/32)2^{-n/2+1}+2^{-(n-1)}$, whereas the linear bias of the functions constructed in Corollary 2 is $2^{-(n-2)/2}$ which is almost optimal, the algebraic immunity is at least $\lceil(n-2)/4\rceil$, and hence the fast algebraic immunity is at least $1+\lceil(n-2)/4\rceil$. So for 1-resilient functions, the functions constructed in [24] and those in Corollary 2 represent two distinct trade-off points with respect to algebraic immunity, fast algebraic immunity and linear bias. From an implementation point of view, however, there is a major difference between the constructions in Corollary 2 and that in [24]. The functions constructed using Corollary 2 require a circuit size of $O(n)$ gates, while the construction in [24] is based on defining the support of the desired Boolean function using powers of a primitive element over the field $\mathbb{F}_{2^{n/2}}$; in particular, from the description of Construction 2 in [24] it appears that the only reasonable method of implementing the function is to use a truth table requiring $\Omega(2^{n})$ gates. We provide a concrete example to highlight the difference between the functions constructed using Corollary 2 and Theorem 6 of [24] (which is based on Construction 2 of [24]). Suppose $n=32$. Theorem 6 of [24] provides a 32-variable, 1-resilient function (say $g_{32}$) with algebraic immunity 16, fast algebraic immunity at least 26 and nonlinearity at least 2147192232 (equivalently, linear bias at most about $2^{-12.85}$) which can be implemented using around $2^{32}$ gates. In Corollary 2 if we take $n=102$, then we obtain a 102-variable, 1-resilient function (say $g_{102}$) with algebraic immunity at least 25, and hence fast algebraic immunity at least 26, and almost optimal linear bias equal to $2^{-50}$ which can be implemented using 406 NAND, 52 XOR, 52 AND, and 3 OR gates (see [7] for the method of obtaining the gate count). So $g_{102}$ and $g_{32}$ both are 1-resilient and have the same fast algebraic immunity, $g_{102}$ has much higher algebraic immunity and much lower linear bias than $g_{32}$; and $g_{102}$ can be implemented much more efficiently than $g_{32}$. From both security and efficiency points of view, to a designer of cryptosystesm $g_{102}$ will be much more preferable than $g_{28}$.
@@
The function $f_{5}$ is a 5-variable, 1-resilient function having degree 3, algebraic immunity 2, nonlinearity 12 (and hence almost optimal linear bias equal to $2^{-2}$), and can be implemented using 7 XOR and 4 AND gates.
@@
This shows that $f_{5}$ is the concatenation of 4 linear functions each of which is non-degenerate on at least 2 variables. It follows that $f_{5}$ is 1-resilient. It is easy to see that the degree of $f_{5}$ is 3. Further, it is not difficult to verify that the distance of $f_{5}$ to any affine is one of the values 12, 16 or 20, and so the nonlinearity of $f_{5}$ is 12. Hence, $f_{5}$ has almost optimal linear bias equal to $2^{-2}$. Since the degree of $f_{5}$ is 3, its algebraic immunity is at most 3. It is easy to see that neither $f_{5}$ nor $1\oplus f_{5}$ has any annihilator of degree 1. The following function is an annihilator of $f_{5}$: $Z_{1}X_{1}\oplus Z_{1}+Z_{2}\oplus X_{2}\oplus Z_{2}\oplus Z_{3}\oplus X_{1}\oplus Z_{3}X_{2}\oplus X_{1}X_{2}\oplus 1$. So the algebraic immunity of $f_{5}$ is 2. The expression for $f_{5}$ given by (6) can be implemented using 7 XOR and 4 AND gates.
@@ 5 Achieving Provable Resiliency/Nonlinearity/Algebraic Immunity Trade-Offs
Any $m_{0}$-resilient function is non-degenerate on at least $m_{0}+1$ variables, any function with linear bias at most $2^{-x_{0}}$ is non-degenerate on at least $2x_{0}$ variables, and any function with algebraic immunity at least $a_{0}$ is non-degenerate on at least $2a_{0}-1$ variables. The first statement now follows from Proposition 1. To see the second statement, note that in all the four cases of Theorem 9, the lower bound on $n$ is linear in $\max(m_{0},x_{0},a_{0})$.